Cystic fibrosis is an autosomal recessive disease caused by a mutation in a single gene with two alleles. Suppose the frequency of cystic fibrosis in a population is 3%. Assuming the gene is in HW equilibrium, calculate the allele and genotype frequencies. If the population consists of 100 individuals, how many are carriers (but unaffected) for the cystic fibrosis allele?
Answer #1

Answer:

According to the question frequency of recessive genotype /cystic fibrosis in the population (q^2)=3/100

So frequency of recessive allele(q)=√0.03=0.17

According to HW equation p+q=1

So p=1-q =0.83( frequency of dominant allele)

Frequency of dominant gene(p^2)=(0.83)^2=0.6889~0.69

Frequency of carriers (2pq)=2×0.83×0.17=0.28

To cross check we can add p^2+2pq+q^2 which should be 1

0.69+0.28+0.03=1

So in a population of 100 there are 28 carriers.

Give a thumbs up if you like the answer. Thank you